please help me solve this question....it says:
a sequance is a special set of numbers. for each of the sequences below:
A) write in words a rule to get each number from its number position in the sequence.
B) write the rule in B in symbols, using n for the number position in the sequence.
the numbers are: a) 3,5,7,9,11 and b) 20,17,14,11

Answer by theflu(4)

A) a)add 2 each time to get the next number
b)subract 3 each time to get the next number
B) a)tn = a + (n - 1)d
in your case:
a = 3 ("a" means first number)
d = 2 ("d" means common difference)
So your formula becomes: tn = 3 + (n - 1)(2)
simplify: tn = 3 + 2n - 2
collect liketerms: tn = 2n + 1 <==answer
to test the answer: try the fifth number, which is suppose to be 11:
t5 = 2(5) + 1 = 11
B) b)try yourself (hint: a = 20; d = -3)